# Quantum transistors for heat flux in and out of working substance parts: harmonic vs transmon and Kerr environs

### v1

@@ Abstract
Quantum thermal transistors have been widely studied in the context of three-qubit systems, where each qubit interacts separately with a Markovian harmonic bath. Markovianity is an assumption that is imposed on a system if the environment loses its memory within short while, while non-Markovianity is a general feature, inherently present in a large fraction of realistic scenarios. Instead of Markovian environments, here we propose a transistor in which the interaction between the working substance and an environment comprising of an infinite chain of qutrits is based on periodic collisions. We refer to the device as a working-substance thermal transistor, since the model focuses on heat currents flowing in and out of each individual qubit of the working substance to and from different parts of the system and environment. We find that the transistor effect prevails in this apparatus and we depict how the amplification of heat currents depends on the temperature of the modulating environment, the system-environment coupling strength and the interaction time. We further show that there exists a non-zero amplification even if one of the environments, that is not the modulating one, is detached from the system. Additionally, the environment, being comprised of three-level systems, allows us to consider the effects of frail perturbations in the energy-spacings of the qutrit, leading to a non-linearity in the environment. We consider non-linearities that are either of transmon- or of Kerr-type. We find parameter ranges where there is a significant amplification for both transmon- and Kerr-type non-linearities in the environment. Finally, we detect the non-Markovianity induced in the system from a non-monotonic behavior of the amplification observed with respect to time, and quantify it using the distinguishability-based measure of non-Markovianity.
@@ I Introduction
Current QTT models typically utilize a three-qubit working substance, each connected to a Markovian bath. Other models include qubit-qutrit systems Guo et al. (2018) and those employing periodic control, such as the Floquet quantum thermal transistor Gupt et al. (2022). There are extensive works which consider Coulomb-coupled conductors Yang et al. (2019); Zhang et al. (2018), polarons in the context of non-equilibrium systems Wang and Xu (2020), double cavity optomechanical systems Xiong et al. (2018), and three terminal normal-superconductors Tang et al. (2019) as models for QTTs. QTT performance, including environmental effects with and without noise, has been extensively studied in single and multi-transistor systems with shared baths. Ekanayake et al. (2023a); Liu et al. (2021); Ekanayake et al. (2023b); Mandarino (2022). The significance of quantum coherence in the context of QTT has also been studied, which enables to provide quantum advantage in the amplification of a QTT over its classical counterpart Su et al. (2021). Further studies on QTT focus on negative differential thermal conductance (NDTC) for high amplification rates and the role of strong coupling and quantum coherence in achieving NDTC Wang et al. (2018); Liu et al. (2019); Wu et al. (2024); Su et al. (2018). In cases requiring high amplification, transistors are often arranged in a Darlington configuration, where the output of one transistor connects to the input of another, greatly increasing the amplification factor. Models similar to these classical ones, but utilizing QTTs have also been rigorously explored. There are instances where such QTTs are placed in different configurations like the Darlington configuration Wijesekara et al. (2021, 2022) or a network of transistors, such as in Craven and Nitzan (2017), to attain many-fold advantage over scenarios that do not involve them. Experimental realization of QTTs using superconducting circuits Gubaydullin et al. (2022); Majland et al. (2020), materials like $VO_{2}$ Ordonez-Miranda et al. (2016); Prod’Homme et al. (2017), and implementation using magnetic thermal transistor Castelli et al. (2023) have been carried out.
@@
Previous QTT studies utilize a global heat current approach, treating the system as three qubits where the focus is on the heat flowing from each bath into the entire working substance. The heat current is ”global” in a sense that the heat current corresponding to each of the baths flow globally into the entire working substance. This paper focuses on local heat currents, where we concentrate on the heat currents flowing “locally” into each qubit from different parts of the system and environment, in contrast to the global heat current approach. We have considered this approach since we know the characteristics of the sub-parts of the working substance experimentally, and it is essential to control and regulate the heat flow in the sub-parts of the system in order to preserve the material for actual implementations in circuits. Our concept of such local heat current is novel, and is hitherto unexplored in literature. There have been previous works which discuss the different types of heat currents in thermal equilibrium Hofer et al. (2017), where the system is connected to an Ohmic bath with a given spectral density. In Ref. Hewgill et al. (2021), for instance, it has been rigorously proven how local master equations are consistent with the laws of thermodynamics. As Markovian nature of a thermal environment, in general, arises under some specific assumptions, and is very unlikely to obtain in a practical scenario, it is important to study the effect of non-Markovianity on transistor action. Quantum thermal devices in the context of non-Markovian environments have been previously delved into Fischer and Breuer (2007); Bhattacharya et al. (2017); Bhattacharyya et al. (2025), which mainly focused on spin-environments. However, here we consider the system-environment interaction to be modeled by a collisional model comprising an infinite array of qutrits, which is experimentally feasible to implement and is fundamentally different from the Markovian environment scenario.
@@
We provide a prototype of a QTT, which consists of three qubits interacting with each other and each qubit is connected separately to an environment. The interaction between the system and environment mimics the collisional model Bäumer et al. (2019); Scarani et al. (2002); Bruneau et al. (2014); Grimmer et al. (2016); Strasberg et al. (2017), which has been extensively studied in literature in the context of quantum batteries Sen and Sen (2023); Morrone et al. (2023), equilibrium and non-equilibrium dynamics Scarani et al. (2002); Bruneau et al. (2014); Grimmer et al. (2016); Seah et al. (2019); Cattaneo et al. (2021), strong coupling thermodynamics Strasberg (2019) and the impact of quantum coherence in thermodynamics Rodrigues et al. (2019); Hammam et al. (2021). In the model that we consider, each system qubit sequentially interacts with a qutrit of the environment for a short but fixed duration of time. The environment, which originally consists of infinite number three-level systems with equispaced energy levels, also allows us to consider the effects of frail perturbations in the energies of the environment-particles, leading to unequal energy spacings of each qutrit. Considering this model as the environment instead of the usual Markovian bosonic one, introduces an inherent non-Markovinity in the system. Using this model, we devise our transistor, which can control and regulate heat currents in the sub-parts of the working substance.
@@
In this work, we scrutinize a class of transistors comprising a three-qubit working substance which interacts with the environment via the collisional model. We find significant amplification of the heat currents at two terminals of the transistor, being modulated by the heat current at the remaining terminal. The amplification factor depends on the temperature of the modulating environment, and system-environment coupling. Specifically, the amplification factors for the left and right qubits first increase and then decrease, as the modulating temperature rises. The amplification trend with coupling strength is complex: it decreases below a critical temperature but increases beyond it with stronger interactions. We also observe some amplification with fewer than three local environments. Additionally, in a linear system with equispaced energy levels, small fluctuations may occur due to external perturbations. Two examples of where such a situation naturally arises are transmons Koch et al. (2007); Josephson (1962); Ray et al. (2023a); Krause et al. (2022); Bhattacharyya et al. (2024) and Kerr-type environments Ray et al. (2023b). In our work, apart from considering non-Markovian environment, which constitutes the collisional model of qubits interacting sequentially with equispaced qutrits, we also consider the effects of inherent non-linearities in the environment such as in transmon and Kerr-types, in the amplification of a type of transistor. We find that amplification and its temperature dependence persist even with small non-linearities in the qutrit energy levels. Transmon and Kerr-type non-linearities enhance amplification in different temperature regimes. Below a critical temperature, Kerr non-linearities outperform linear qutrits, while above it, transmon qutrits are more effective. Furthermore, we find a non-monotonic behavior in the amplification factor of the three-qubit transistor setup with respect to time, and quantify the non-Markovianity that possibly induces such behavior.
@@
The paper is organized as follows. In Sec. II, we explicitly describe the collisional model for the system and environment that we consider, along with defining the relevant Hamiltonians, the dynamical amplification factor, and the different types of heat-currents. We then analyze the effect of varying temperature of the modulating bath, and coupling constant between the bath and the working substance on the amplification considering linear baths in Sec. III. We consider both the cases where there are three (Sec. III.1) and less than three (Sec. III.2) environments. We then introduce non-linearity in the baths and analyze its effect on amplification in Sec. IV. In Sec. V, the non-Markovianity present in the system is quantified and analyzed. Finally the concluding remarks are presented in Sec. VI.
@@
Figure 1: Schematic representation of the difference between the local and global approaches utilized to provide the definition of heat current used in previous models, referred to as BTTs versus the model studied in this paper, referred to as WTTs. The name, BTT essentially means that the transistor is conceptualized by looking at heat current flows in and out of each of the local baths to and from the entire system, as depicted by the red arrows in panel (a), and hence the transistor action is scrutinized by focusing on the “bath”. The name, WTT suggests that, the heat is exchanged between each (local) qubit with the different parts of the system and the environment, as shown by the blue arrows in panel (b), and therefore the focus is on the “working substance”. BTTs correspond to the usual QTT model previously studied in literature, while the class of transistor devices which we analyze in this paper are the WTTs.
@@ II.2 Relevant Hamiltonians
$H_{sys}=-\sum_{i\in\{L,M,R\}}\frac{\hbar\omega_{i}}{2}\leavevmode\nobreak\ \sigma_{Z}^{i}-\sum_{\begin{subarray}{c}i,j\in\{L,M,R\}\\ i\neq j\end{subarray}}\frac{\hbar\omega_{ij}}{2}\leavevmode\nobreak\ \sigma_{Z}^{i}\leavevmode\nobreak\ \sigma_{Z}^{j},$ (2)
@@
$H_{env}=-\hbar\leavevmode\nobreak\ \Delta\sum_{i\in\{L,M,R\}}\sigma_{Z}^{(i),env},$ (3)
@@
In Sec. III, we consider the collisional model comprising of equispaced qutrits, which arrive one at a time to interact with the system qubit. We refer to this type of environment as linear because of the equal spacing between the energy levels of the three-level systems. Each of the three qubits interact locally with each qutrit of the environment via the qubit-qutrit interaction given by
$H_{qubit-env}=-\hbar\leavevmode\nobreak\ g\sum_{i\in\{L,M,R\}}\sigma_{X}^{(i),qubit}\leavevmode\nobreak\ \sigma_{X}^{(i),env}$ (4)
where $\sigma_{X}^{(i),qubit}$ is the Pauli-$x$ matrix for spin-$1/2$ particle, $\sigma_{X}^{(i),env}$ is the Pauli-$x$ matrix for spin-$1$ particle, and $g$ is the coupling constant between the system-qubit and the environment-qutrit. The quantities $\omega_{i},\omega_{ij},\Delta$ and $g_{k}$ are all in units of time-1, $\forall\leavevmode\nobreak\ i,j$.
For simplicity we consider $\omega_{L}=\omega_{M}=\omega_{R}=\omega_{ML}=\omega_{MR}=\Delta$ and $\omega_{LR}=0$. The disparity between the values of the interaction strengths $\omega_{ML}=\omega_{MR}$ and $\omega_{LR}$ introduces asymmetry in the system, which ensures transistor effect to be present.

### v2

@@ Abstract
Quantum thermal transistors have been widely studied in the context of three-qubit systems, where each qubit interacts separately with a Markovian harmonic bath. Markovianity is an assumption that is imposed on a system if the environment loses its memory within short while, while non-Markovianity is a general feature, inherently present in a large fraction of realistic scenarios. Instead of Markovian environments, here we propose a transistor in which the interaction between the working substance and an environment comprising of an infinite chain of qutrits is based on periodic collisions. We refer to the device as a working-substance thermal transistor, since the model focuses on heat currents flowing in and out of each individual qubit of the working substance to and from different parts of the system and environment. We find that the transistor effect prevails in this apparatus and we depict how the amplification of heat currents depends on the temperature of the modulating environment, the system-environment coupling strength and the interaction time. Initially we consider no direct coupling between the left and right qubits, and only the left-middle and right-middle terminals are coupled. Later, we also investigate both symmetric and asymmetric coupling configurations where the left and right qubits are coupled to each other, and observe amplification in the left and right qubits in both the cases. We further show that there exists a non-zero amplification even if one of the environments, that is not the modulating one, is detached from the system. Additionally, the environment, being comprised of three-level systems, allows us to consider the effects of frail perturbations in the energy-spacings of the qutrit, leading to a non-linearity in the environment. We consider non-linearities that are either of transmon- or of Kerr-type. We find parameter ranges where there is a significant amplification for both transmon- and Kerr-type non-linearities in the environment. The analysis is performed under the condition of vanishing left-right coupling, along with the symmetric and asymmetric coupling configurations as well. Furthermore, we demonstrate a prominent transistor effect for the working substance when coupled to qubit environments, within the same collisional model framework. Finally, we detect the non-Markovianity induced in the system from a non-monotonic behavior of the amplification observed with respect to time, and quantify it using the distinguishability-based measure of non-Markovianity, for all the three cases mentioned above.
@@ I Introduction
Current QTT models typically utilize a three-qubit working substance, each connected to a Markovian bath. Other models include qubit-qutrit systems Guo et al. (2018) and those employing periodic control, such as the Floquet quantum thermal transistor Gupt et al. (2022); Das et al. . There are extensive works which consider Coulomb-coupled conductors Yang et al. (2019); Zhang et al. (2018), polarons in the context of non-equilibrium systems Wang and Xu (2020), double cavity optomechanical systems Xiong et al. (2018), and three terminal normal-superconductors Tang et al. (2019) as models for QTTs. QTT performance, including environmental effects with and without noise, has been extensively studied in single and multi-transistor systems with shared baths. Ekanayake et al. (2023a); Liu et al. (2021); Ekanayake et al. (2023b); Mandarino (2022). The significance of quantum coherence in the context of QTT has also been studied, which enables to provide quantum advantage in the amplification of a QTT over its classical counterpart Su et al. (2021). Further studies on QTT focus on negative differential thermal conductance (NDTC) for high amplification rates and the role of strong coupling and quantum coherence in achieving NDTC Wang et al. (2018); Liu et al. (2019); Wu et al. (2024); Su et al. (2018). Models analogous to these classical ones, but employing QTTs, have also been rigorously investigated. In particular, configurations such as the Darlington setup Wijesekara et al. (2021, 2022), where the output of one transistor serves as the input to another, or more complex transistor networks as considered in Craven and Nitzan (2017), have been shown to provide significant performance enhancements compared to scenarios that do not incorporate QTTs. Experimental realization of QTTs using superconducting circuits Gubaydullin et al. (2022); Majland et al. (2020), materials like $VO_{2}$ Ordonez-Miranda et al. (2016); Prod’Homme et al. (2017), and implementation using magnetic thermal transistor Castelli et al. (2023) have been carried out.
@@
Previous QTT studies utilize a global heat current approach, treating the system as three qubits where the focus is on the heat flowing from each bath into the entire working substance. The heat current is ”global” in a sense that the heat current corresponding to each of the baths flow globally into the entire working substance. This paper focuses on local heat currents, where we concentrate on the heat currents flowing “locally” into each qubit from different parts of the system and environment, in contrast to the global heat current approach. We have considered this approach since we know the characteristics of the sub-parts of the working substance experimentally, and it is essential to control and regulate the heat flow in the sub-parts of the system in order to preserve the material for actual implementations in circuits. Our concept of such local heat current is novel, and is hitherto unexplored in literature. There have been previous works which discuss the different types of heat currents in thermal equilibrium Hofer et al. (2017), where the system is connected to an Ohmic bath with a given spectral density. Here, they have used certain approximations to derive the local master equation and subsequently the definition of the local heat current. These assumptions are as follows. Firstly, the interaction between system and environment is perturbatively very small. Secondly, the Markov approximation assumption is made which tells about the invariance of system’s state on time scales of order correlation time. However, these assumptions are not valid in our case. We have defined our system for arbitrary coupling instead of weak coupling which does not follow the Born approximation. The Markov approximation also talks about the bath retaining its state during and after the evolution, which is not true in the dynamics we have considered. Our definition of a local heat current is not equivalent to their definition since the Born Markov assumptions on which they have defined the local heat current are not valid in our system.
In our case, the working substance is considered under the most general completely positive trace-preserving (CPTP) dynamics, i.e. its evolution is described by a global unitary acting on the joint system of the working substance and its environment, followed by discarding the environment. Importantly, no assumption of Markovianity is made in the analysis, and therefore, Markovian master equations do not naturally arise in this framework.
In Ref. Hewgill et al. (2021), for instance, it has been rigorously proven how local master equations are consistent with the laws of thermodynamics. As Markovian nature of a thermal environment, in general, arises under some specific assumptions, and is very unlikely to obtain in a practical scenario, it is important to study the effect of non-Markovianity on transistor action. Quantum thermal devices in the context of non-Markovian environments have been previously delved into Fischer and Breuer (2007); Bhattacharya et al. (2017); Bhattacharyya et al. (2025); Mondkar et al. (2025), which mainly focused on spin-environments. However, here we consider the system-environment interaction to be modeled by a collisional model comprising an infinite array of qutrits, which is experimentally feasible to implement and is fundamentally different from the Markovian environment scenario.
@@
We provide a prototype of a QTT, which consists of three qubits interacting with each other and each qubit is connected separately to an environment. The interaction between the system and environment mimics the collisional model Bäumer et al. (2019); Scarani et al. (2002); Bruneau et al. (2014); Grimmer et al. (2016); Strasberg et al. (2017), which has been extensively studied in literature in the context of quantum batteries Sen and Sen (2023); Morrone et al. (2023), equilibrium and non-equilibrium dynamics Scarani et al. (2002); Bruneau et al. (2014); Grimmer et al. (2016); Seah et al. (2019); Cattaneo et al. (2021), strong coupling thermodynamics Strasberg (2019) and the impact of quantum coherence in thermodynamics Rodrigues et al. (2019); Hammam et al. (2021). In the model that we consider, each system qubit sequentially interacts with a qutrit of the environment for a short but fixed duration of time. The environment, which originally consists of infinite number three-level systems with equispaced energy levels, also allows us to consider the effects of frail perturbations in the energies of the environment-particles, leading to unequal energy spacings of each qutrit. Considering this model as the environment instead of the usual Markovian bosonic one, introduces an inherent non-Markovianity in the system. Using this model, we devise our transistor, which can control and regulate heat currents in the sub-parts of the working substance.
@@
In this work, we scrutinize a class of transistors comprising a three-qubit working substance which interacts with the environment via the collisional model. There also exist interactions between the components of the working substance, i.e. the left, middle, and right qubits. Initially, we consider equal interaction strength between the left-middle and right-middle qubits, but the left and right qubits remain non-interacting. We find significant amplification of the heat currents at two terminals of the transistor, being modulated by the heat current at the remaining terminal. The amplification factor depends on the temperature of the modulating environment, and system-environment coupling. Specifically, the amplification factors for the left and right qubits first increase and then decrease, as the modulating temperature rises. The amplification trend with coupling strength is complex: it decreases below a critical temperature but increases beyond it with stronger interactions. We also observe some amplification with fewer than three local environments. Additionally, in a linear system with equispaced energy levels, external perturbations can induce small fluctuations, effectively rendering the system nonlinear. Two examples of where such a situation naturally arises are transmons Koch et al. (2007); Josephson (1962); Ray et al. (2023a); Krause et al. (2022); Bhattacharyya et al. (2024) and Kerr-type environments Ray et al. (2023b). In our work, apart from considering non-Markovian environment, which constitutes the collisional model of qubits interacting sequentially with equispaced qutrits, we also consider the effects of non-linearities in the environment such as in transmon and Kerr-types, in the amplification of a type of transistor. We find that amplification and its temperature dependence persist even with small non-linearities in the qutrit energy levels. Transmon and Kerr-type non-linearities enhance amplification in different temperature regimes. Below a critical temperature, Kerr non-linearities outperform linear qutrits, while above it, transmon qutrits are more effective. Furthermore, we find a non-monotonic behavior in the amplification factor of the three-qubit transistor setup with respect to time, and quantify the non-Markovianity that possibly induces such behavior. Furthermore, we consider two cases, i.e. symmetric and asymmetric coupling, where for symmetric coupling, left-middle, middle-left, and left-right qubits of working substance are coupled with each other by equal interaction strength. On the other hand, for asymmetric coupling, left-middle, middle-left, and left-right qubits are coupled by different interaction strengths. We show that appreciable amplification is obtainable for both the cases corresponding to both left and right qubits. Moreover, we also observe the benefits of nonlinear environments in obtaining higher amplification in both symmetric and asymmetric cases. Finally, instead of considering the environment to be a collision model composed of qutrits, we consider the collision model to consist of a qubit system and again show sufficient amplification corresponding to both right and left qubits. Moreover, we analyze the effects of non-linearities present in the environment in this setting. We scrutinize the distance-based non-Markovianity measure in this situation as well.
@@
The paper is organized as follows. In Sec. II, we explicitly describe the collisional model for the system and environment that we consider, along with defining the relevant Hamiltonians, the dynamical amplification factor, and the different types of heat-currents. We then analyze the effect of varying temperature of the modulating bath, and coupling constant between the bath and the working substance on the amplification, along with the symmetric and asymmetric cases, considering linear baths in Sec. III. We consider both the cases where there are three (Sec. III.1) and less than three (Sec. III.2) environments. We then introduce non-linearity in the baths and analyze its effect on amplification in Sec. IV. In Sec. V, the non-Markovianity present in the system is quantified and analyzed. Analysis of the transistor effect using qubits as components of the environment is provided in Sec. VI. Finally the concluding remarks are presented in Sec. VII.
@@
Figure 1: Schematic representation of the difference between the local and global approaches utilized to provide the definition of heat current used in previous models, referred to as BTTs versus the model studied in this paper, referred to as WTTs. The name, BTT essentially means that the transistor is conceptualized by looking at heat current flows in and out of each of the local baths to and from the entire system, as depicted by the red arrows in panel (a), and hence the transistor action is scrutinized by focusing on the “bath”. The name, WTT suggests that, the heat is exchanged between each (local) qubit with the different parts of the system and the environment, as shown by the blue arrows in panel (b), and therefore the focus is on the “working substance”. BTTs correspond to the usual QTT model previously studied in literature, while the class of transistor devices which we analyze in this paper are the WTTs. The bath interacts with the working substance via the collisional model where every qutrit sequentially interacts with the working substance. We show in the schematic that the $n^{\text{th}}$ qubit is interacting with the system while the $(n-1)^{\text{th}}$ qubit has already interacted. The interaction with the $(n+1)^{\text{th}}$ qubit is going to occur.
@@ II.2 Relevant Hamiltonians
$H_{sys}=-\sum_{i\in\{L,M,R\}}\frac{\hbar\omega_{i}}{2}\penalty 10000\ \sigma_{z}^{i}-\sum_{\begin{subarray}{c}i,j\in\{L,M,R\}\\ i\neq j\end{subarray}}\frac{\hbar\omega_{ij}}{2}\penalty 10000\ \sigma_{z}^{i}\penalty 10000\ \sigma_{z}^{j},$ (2)
@@
$H_{env}=-\hbar\penalty 10000\ \Delta\sum_{i\in\{L,M,R\}}\sigma_{z}^{(i),env},$ (3)
@@
In Sec. III, we consider the collisional model comprising of equispaced qutrits, which arrive one at a time to interact with the system qubit. Each of the three qubits interact locally with each qutrit of the environment via the qubit-qutrit interaction given by
$H_{qubit-env}=-\hbar\penalty 10000\ g\sum_{i\in\{L,M,R\}}\sigma_{x}^{(i),qubit}\penalty 10000\ \sigma_{x}^{(i),env}$ (4)
where $\sigma_{x}^{(i),qubit}$ is the Pauli-$x$ matrix for spin-$1/2$ particle, $\sigma_{x}^{(i),env}$ is the Pauli-$x$ matrix for spin-$1$ particle, and $g$ is the coupling constant between the system-qubit and the environment-qutrit. The quantities $\omega_{i},\omega_{ij},\Delta$ and $g_{k}$ are all in units of time-1, $\forall\penalty 10000\ i,j$. For simplicity we consider $\omega_{L}=\omega_{M}=\omega_{R}=\omega_{ML}=\omega_{MR}=\Delta$ and $\omega_{LR}=0$, unless stated otherwise.
The environment is modeled as an infinite sequence of three-level systems or qutrits, each of which sequentially interacts with a single qubit comprising the working substance. Each qutrit interacts with the qubit for a fixed duration $\delta t$, after which it is replaced by a new, identical qutrit. As a result, the environment effectively resets after each interaction interval $\delta t$. Due to this periodic interaction between the system and environmental units, this framework is commonly referred to in the literature as a collisional model. The dynamical nature of the environment - whether Markovian or non-Markovian - is determined by key parameters of the model, such as the coupling strength between the system and the environment and the collision time. If the interaction time between the system and each environmental qutrit is sufficiently small for a given interaction strength, the environment is able to extract information from the system but lacks the time necessary to return it. This results in Markovian dynamics, characterized by a lack of memory effects. Conversely, when the interaction time with each qutrit is increased, the environment can not only retain information about the system but also feed it back during subsequent interactions. This backflow of information leads to non-Markovian dynamics, wherein memory effects play a significant role. On the other hand, if the environmental qutrits interact perturbatively with the system, i.e., the interaction strength is very weak, and are subsequently replaced after each interaction, as prescribed by the collisional model, the environment fails to retain sufficient information to influence the evolution of the system. As a result, the system exhibits Markovian dynamics, characterized by the absence of memory effects. In contrast, when the system–environment interaction strength is increased, the environment becomes more entangled with the system and retains information about its past states. If the structure of subsequent interactions permits, this stored information can flow back into the system, thereby influencing its future dynamics and giving rise to non-Markovian behavior.